OpenStudy (anonymous):

The polynomial x^2 -4x +3 is a factor of x^2 +(a-4)x^2 + (3-4a)x +3 calculate the value of a constant a

OpenStudy (anonymous):

hmm i guess 1 is a zeros of the first quadratic, so it must be a zero of the second one as well. replace x by 1, set the result equal to zero, and solve for a

OpenStudy (anonymous):

i have an answer i think is right. what do you get?

OpenStudy (anonymous):

i dont get it hahah

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

\[x^2 +(a-4)x^2 + (3-4a)x +3 =( x^2 -4x +3)(\text{ something})\]

OpenStudy (anonymous):

if you let \[x=1\] you get on the right hand side \[(1-4+3)(\text{something})=0\]

OpenStudy (anonymous):

which means (since they are equal) the left hand side must be zero as well if you let x = 1

OpenStudy (anonymous):

so replace x on the left hand side by 1, set the result equal zero, and you can solve for a

OpenStudy (anonymous):

you with me so far?

OpenStudy (anonymous):

ahhhhh ok ! yes am with you!

OpenStudy (anonymous):

is it just sorta like trial and era?

OpenStudy (anonymous):

ok so now we replace x by one and get \[1+(a-4)+ (3-4a) +3 =0\] \[3-3a=0\] \[a=1\]

OpenStudy (anonymous):

no not really if we know that 1 is a zero of \[x^2-4x+3\] then since this is a factor of the other polynomial, it must be a zero of that one as well

OpenStudy (anonymous):

ahhhh ok!! :D

OpenStudy (anonymous):

and could it be anything else?

OpenStudy (anonymous):

like x=3

OpenStudy (anonymous):

notice that if we had just plugged in any number we would not be in such good shape. it is the zero that makes this work.

OpenStudy (anonymous):

maybe. lets try it

OpenStudy (anonymous):

\[x^2-4x+3=(x-3)(x-1)\] so 3 is a zero as well. so we can replace x by 3 on the left hand side and see what we get \[3^2+(z-4)\times 3^2+(3-4a)\times 3+3=0\]

OpenStudy (anonymous):

hmm looks like i get \[a=-5\]

OpenStudy (anonymous):

man i must have messed this up, or else there is no solution. in fact is \[x^2-4x+3\] is a factor, then both 1 and 3 must give zero. but you get different values for a depending on whether you use 1 or 3, so i am going to say there is no solution to this

OpenStudy (anonymous):

you sure it is \[x^2 +(a-4)x^2 + (3-4a)x +3 \]

OpenStudy (anonymous):

yapp am sure it is that!!

OpenStudy (anonymous):

hmm but then in the first one shoudnt it be plus 4 in the end and not 3? so its 1+(a-4) + (3-4a) +4?

OpenStudy (anonymous):

ok so i am thinking there is no solution to this one if \[x^2 +(a-4)x^2 + (3-4a)x +3 =(x-1)(x-3)(\text{something})\] then that means they are equal for all values of x. now if x = 1 i get a =1 and if x = 3 i get z = -5, and since they are not the same, there is a big problem here

OpenStudy (anonymous):

no, if x = 1 you should get \[1+(a-4)+(3-4a)+3=0\]

OpenStudy (anonymous):

ohh ok hahaha

OpenStudy (anonymous):

hmm then i guess there is a huge problem

OpenStudy (anonymous):

i have another problem as well. the constant on both sides is 3

OpenStudy (anonymous):

really?? :O

OpenStudy (anonymous):

oh my this question is bad retrice

OpenStudy (anonymous):

and both sides are of degree 2 so if the left hand side is a multiple of the right hand side, then it is a constant multiple

OpenStudy (anonymous):

yes i see your point

OpenStudy (anonymous):

and since they both have constant 3, that constant is 1

OpenStudy (anonymous):

in other words if one is a multiple of the other, they are equal

OpenStudy (anonymous):

YAH YAH

OpenStudy (anonymous):

is this question from a book of some kind?

OpenStudy (anonymous):

nope my teacher uploaded it! it says exercises from may examinations 2004 haha

OpenStudy (anonymous):

i got a whole pellet to do.. :O plus questions in the book

OpenStudy (anonymous):

maybe there is a typo

OpenStudy (anonymous):

because look, if the sides are equal, that means \[x^2 +(a-4)x^2 + (3-4a)x +3=x^2-4x+3\] which means \[a-4=0\] or \[a=4\] but if \[a=4\]then \[3-4a=-13\] which is impossible since it has to be -4 what class is this? i really think there is a mistake here can you upload the question directly?

OpenStudy (anonymous):

if you send me the link i will look more carefully, because i would hate to think that i am telling you something that is wrong. but at this point i would say there is a mistake

OpenStudy (anonymous):

ok its not a link though its a pdf file..

OpenStudy (anonymous):

ok attach it and i will take a look